OpenStudy (darkigloo):

calculus The center of mass of quarter circle y=sqrt( r^2 -x^2) is given by the point??

OpenStudy (darkigloo):

\[y= \sqrt{r^2 - x^2} , x \in[0,r]\]

OpenStudy (bobo-i-bo):

Given a lamina, do you know the formula to calculate the centre of mass? (Do you know what a lamina is?)

OpenStudy (darkigloo):

i dont know what a lamina is

OpenStudy (bobo-i-bo):

A lamina is a flat shape with constant density. in this case the lamina is the quarter circle.

OpenStudy (darkigloo):

ok.

OpenStudy (bobo-i-bo):

There are one or two formulae for finding the centre of mass of a lamina, but the one you want is: \[\frac{\int\limits_{a}^{b}xydx}{\int\limits_{a}^{b}ydx}\] which will give you the centre of mass

OpenStudy (bobo-i-bo):

Understand what to do from here? It would be a very good idea if you tried to search up where the formula comes from though btw

OpenStudy (darkigloo):

\[\frac{ \int\limits_{0}^{r} x(\sqrt{r ^{2}-x ^{2}})dx }{ \int\limits_{0}^{r} \sqrt{r ^{2}-x ^{2}}dx } \]

OpenStudy (darkigloo):

like that?

OpenStudy (bobo-i-bo):

yup!

OpenStudy (darkigloo):

how would i find that integral?

OpenStudy (mathmale):

the integral in the numerator can be solved using a substitution. The one in the denom. can be solved (integrated) using a trig substitution.

OpenStudy (darkigloo):

can you help me with that? \[\sqrt{a^2 - x^2} \rightarrow x=a \sin \theta \] \[\sqrt{(asin \theta)^{2}-x^2} \]

OpenStudy (darkigloo):

thats for the denominator. what do i do from there?

OpenStudy (darkigloo):

@mathmale

OpenStudy (darkigloo):

\[x= r \sin \theta \] \[dx = r \cos \theta d \theta \] ?